Calculate the net work done to the environment dur
Answer #1

W = -P*\Delta V

Where

W stands for work done

P stands for pressure

\Delta V for change in volume

l pressure is 10 Pa. 2.0 2.5 5.0

For 1

W = Area under the line which is the sum of area of triangle and rectangle below it

W = 0.5*1.5*105*2.5 + 2*105*2.5 = 187500 + 500000 = 687500 J

Since volume is increasing it will be negative

For 2

It will be the area of the rectangle = 500000J

Since volume is decreasing work will be positive

For 3

Work done is zero as there is no change in volume

Hence net work done

-687500 + 500000 = -187500 J